Recently, More Money 4U offered an annuity that pays 6.3% compounded monthly. If $1,196 is deposited into this annuity every month, how much is in the account after 7 years? How much of this is interest?

Type the amount in the account: $____

Type the amount of interest earned: $____

(Please square, label, or circle the answer)
Answer #1

Amount in the account after 7 years:-

=FV(rate,nper,pmt,pv)

=FV(6.3%/12,7*12,-1196)

=125857.64

Present value of money:-

=PV(6.3%/12,7*12,-1196)

=81069.36

Total amount paid =1196*7*12 =100464

Interest =100464-81069.36 =19394.64